### v2

@@ 1 Introduction
The study of correlations is of central importance in quantum information theory, as it provides a comprehensive understanding of how information is distributed among different parts of a quantum system. Correlations are often explored through various entropy inequalities and related measures, which help in understanding the fundamental limits and capabilities of quantum systems, guiding the development of quantum technologies and deepening our theoretical understanding of quantum mechanics.
Motivated by research in quantum gravity, in particular in the context of gauge/gravity duality Maldacena:1997re ; Gubser:1998bc ; Witten:1998qj , a deep question about fundamental possible “patterns of dependences” in quantum mechanics was raised in Hernandez-Cuenca:2019jpv , which dubbed it the quantum marginal independence problem (QMIP). In its simplest and most intuitive form, the QMIP asks the following question: Given the binary specification of presence versus absence of correlations between all pairs of disjoint subsystems, does there exist any quantum state realizing such a choice? As indicated in the following, the solution to this innocuously simple question is remarkably powerful. Below, we will specify certain aspects of this problem more fully, then motivate and contextualize them, and finally explain our strategy to tackle them.
@@
In quantum gravity, there is strong evidence supporting the conjecture Hernandez-Cuenca:2022pst that this knowledge is in fact sufficient to derive all entropy inequalities that characterize the holographic entropy cone Bao:2015bfa . Specifically, in this context one can define a restricted version of the QMIP, called the holographic marginal independence problem (HMIP), where instead of considering which faces can be realized by arbitrary quantum states, one focuses on the restricted class of “geometric states.”555 These are the states of a holographic CFT that describe a classical bulk geometry, for which the entanglement entropy of a subsystem is computed by the HRRT prescription Ryu:2006bv ; Hubeny:2007xt . For any given number of parties ${\sf N}$, it was argued in Hernandez-Cuenca:2022pst that if one knows the solution to the HMIP for each ${\sf N}^{\prime}\geq{\sf N}$, then one can in principle reconstruct all the inequalities that specify the ${\sf N}$-party holographic entropy cone. In fact, Hernandez-Cuenca:2022pst suggested that there might be an even more remarkable structure which only requires the solution to the “extremal” version of the HMIP rather than the full one. More precisely, let $\mathscr{R}_{{}_{\text{SA}}}^{{}^{{\sf N}}}$ be the set of all $1$-dimensional faces, or extreme rays (ERs), of the SACN, and $\mathscr{R}_{{}_{\text{Hol}}}^{{}^{{\sf N}}}\subseteq\mathscr{R}_{{}_{\text{SA}}}^{{}^{{\sf N}}}$ be the subset of these that can be realized holographically.666 Similarly, we will use $\mathscr{R}^{{}^{{\sf N}}}$ with other subscripts to indicate subsets of the SACN extreme rays pertaining to the class indicated by the subscript; refer to Table 1 for a summary of these. Then Hernandez-Cuenca:2022pst argued that for any ${\sf N}$, there exists some ${\sf N}^{\prime}\geq{\sf N}$ (which depends on ${\sf N}$) such that $\mathscr{R}_{{}_{\text{Hol}}}^{{}^{{\sf N}^{\prime}}}$ provides sufficient information to derive the extreme rays of the ${\sf N}$-party holographic entropy cone via a simple coarse-graining of subsystems. These then allow for a straightforward reconstruction of all the ${\sf N}$-party holographic entropy inequalities (although in practice this might be a hard computation). The result of Hernandez-Cuenca:2022pst is quite striking since the SAC itself is a much more primitive construct than the holographic entropy cone. It originates from the self-evident statement that the amount of correlation between any pair of subsystems cannot be negative, and as such it is not cognizant of holography. Indeed, the holographic input enters through the restriction to $\mathscr{R}_{{}_{\text{Hol}}}^{{}^{{\sf N}}}$.
@@
Interestingly, even though Hernandez-Cuenca:2022pst relied on technology only applicable in the holographic setting, a clue hinting at a similar structure in more general contexts was found recently in He:2023aif , which considered how to derive inner bounds to the general quantum entropy cone 1193790 ; 2005:Linden ; 2011:Cadney ; Christandl_2023 from the solution to the QMIP. Denoting by $\mathscr{R}_{{}_{\text{Sta}}}^{{}^{{\sf N}}}\subseteq\mathscr{R}_{{}_{\text{SA}}}^{{}^{{\sf N}}}$ the set of extreme rays of the SACN that can be realized by stabilizer states, the main result of He:2023aif was that a subset of $\mathscr{R}_{{}_{\text{Sta}}}^{{}^{9}}$ is sufficient to derive777 As in the holographic case, by derive we mean via subsystems coarse-grainings. an inner bound to the $4$-party quantum entropy cone which coincides with the full 4-party stabilizer cone Linden:2013kal .
@@
These results suggest that a similar structure might also exist at a more abstract level. For ${\sf N}\geq 3$, most faces of the SACN cannot be part of the solution to the QMIP for the simple reason that they are “ruled out” by strong subadditivity (SSA). Following He:2022bmi , we will say that a face is SSA-compatible if there is at least one entropy vector in its interior that satisfies all instances of SSA. For a fixed ${\sf N}$, the set of SSA-compatible faces has the structure of a lattice,888 A lattice is a poset $\mathcal{P}$ endowed with the additional property that for any two elements $x,y\in\mathcal{P}$, there exists both a least upper bound (called the join and denoted $x\vee y$) and a greatest lower bound (called the meet and denoted $x\wedge y$). where the partial order is given by reverse inclusion ($\mathscr{F}_{1}<\mathscr{F}_{2}$ if $\mathscr{F}_{1}\supset\mathscr{F}_{2}$), and the meet of two faces is the lowest dimensional face $\mathscr{F}=\mathscr{F}_{1}\wedge\mathscr{F}_{2}$ that includes $\mathscr{F}_{1}$ and $\mathscr{F}_{2}$ on it boundary He:2022bmi .999 The set of faces of an arbitrary polyhedral cone forms a lattice with the same partial order and meet operation. In general, the subset $\mathscr{R}_{{}_{\text{SSA}}}^{{}^{{\sf N}}}\subseteq\mathscr{R}_{{}_{\text{SA}}}^{{}^{{\sf N}}}$ of SSA-compatible extreme rays of the SACN is a subset of the set of coatoms of this lattice,101010 For ${\sf N}\leq 4$, each coatom is an extreme ray, but it is not known if the inclusion is strict for larger ${\sf N}$; see He:2022bmi for more details. but He:2022bmi showed that for ${\sf N}\geq 4$ the lattice is not coatomistic. This means that not every element of the lattice can be obtained as the meet of a collection of coatoms. In other words, even geometrically, knowing $\mathscr{R}_{{}_{\text{SSA}}}^{{}^{{\sf N}}}$ is in general insufficient information to be able to recover all the SSA-compatible higher-dimensional faces.111111 Technically, this result was proven for the set of SSA-compatible faces, which are not necessarily all realizable by quantum states. Therefore, it could have been in principle possible that each element of the lattice which is not the meet of a collection of coatoms is non-realizable, implying that the sublattice of realizable elements is coatomistic. However, we will see momentarily that already for ${\sf N}=4$ this is not true. This result strongly suggests that in general, the solution to the QMIP for fixed ${\sf N}$ cannot be deduced from the solution $\mathscr{R}_{{}_{\text{QM}}}^{{}^{{\sf N}}}$ of its extremal version for the same ${\sf N}$, and that one needs to know all the meet-irreducible elements of the lattice.121212 An element $x$ of a lattice is said to be meet-irreducible if $x=y\wedge z$ implies $x=y$ or $x=z$. For ${\sf N}=4$ however, it was already shown in Hernandez-Cuenca:2019jpv that all SSA-compatible faces of the SAC4 are realizable in quantum mechanics, and that in fact stabilizer states are sufficient. This implies that the meet-irreducible elements of the lattice of SSA-compatible faces of the SAC4 which are not coatoms can be obtained from coarse-grainings of certain coatoms of the analogous lattice for ${\sf N}=9$. In other words, similar to the holographic case, the solution to the QMIP can be derived from the solution to its extremal version involving a larger number of parties. All these observations suggest that there might be classes of quantum states for which the solution to the extremal marginal independence problem for all ${\sf N}^{\prime}\geq{\sf N}$ carries a lot of (if not complete) information about the solution to the (seemingly) more general ${\sf N}$-party QMIP, and even about entropy inequalities. This motivates us to study the set $\mathscr{R}_{{}_{\text{SSA}}}^{{}^{{\sf N}}}$ more closely.
@@
Another important motivation for us to study $\mathscr{R}_{{}_{\text{SSA}}}^{{}^{{\sf N}}}$ comes again from holography. Suppose that the conjecture of Hernandez-Cuenca:2022pst is true, and that all ${\sf N}$-party holographic entropy inequalities can indeed be recovered from $\mathscr{R}_{{}_{\text{Hol}}}^{{}^{{\sf N}^{\prime}}}$ for some ${\sf N}^{\prime}\geq{\sf N}$. Then what characterizes the extreme rays that can be realized in quantum mechanics but not by geometric states? A first guess, informed by data for ${\sf N}\leq 5$, is that no such rays exist and $\mathscr{R}_{{}_{\text{QM}}}^{{}^{{\sf N}}}=\mathscr{R}_{{}_{\text{Hol}}}^{{}^{{\sf N}}}$. In other words, all extreme rays of the SACN that can be realized in quantum mechanics would also be realizable holographically. This however was shown to be false in He:2023cco , which proved the strict inclusion $\mathscr{R}_{{}_{\text{Hol}}}^{{}^{{\sf N}}}\subset\mathscr{R}_{{}_{\text{QM}}}^{{}^{{\sf N}}}$ for ${\sf N}\geq 6$ via the construction of an explicit counterexample, namely an extreme ray of the SAC6 that can be realized by a stabilizer state but not by a geometric state.
@@
where {${n\brace k}$ is the Stirling number of the second kind Hubeny:2018ijt .131313 The number of SA instances is {${{\sf N}+2\brace 3}$, but the $12^{{\sf N}}-$ 1 instances of the form ${\sf S}_{\mathscr{J}}\geq 0$ are not facets He:2022bmi . For ${\sf N}\leq 4$ the problem is easily manageable by freely available software Normaliz:301 , but the ${\sf N}=5$ case is already significantly more involved, since ${\sf{D}}=31$ and there are $270$ facets. The computation can be largely simplified using a result of He:2022bmi , which identified for arbitrary ${\sf N}$ a specific lower-dimensional face $\mathscr{F}_{*}$ of the SACN that includes all elements of $\mathscr{R}_{{}_{\text{SSA}}}^{{}^{{\sf N}}}$ (with the exception of the trivial ones corresponding to the Bell pairs).141414 Since we will utilize this result in the computation presented in this paper, we will review it in detail in the next section. In general, the dimension of $\mathscr{F}_{*}$ is given by
@@
which for ${\sf N}=5$ gives $\text{dim}(\mathscr{F}_{*})=16$. Furthermore, upon dimensional reduction of the inequalities that specify the SAC5 to the linear subspace spanned by $\mathscr{F}_{*}$, many inequalities become redundant and $\mathscr{F}_{*}$ only has $80$ facets. In this manner, the computation of $\mathscr{R}_{{}_{\text{SA}}}^{{}^{5}}$ becomes straightforward using Normaliz:301 , but as one can immediately see from (4) and (5), this simplification quickly becomes insufficient as ${\sf N}$ grows. Indeed, already at ${\sf N}=6$, the face $\mathscr{F}_{*}$ has $\text{dim}(\mathscr{F}_{*})=42$ and is specified by $280$ inequalities (the number of facets of SAC6 is $903$), rendering the computation of $\mathscr{R}_{{}_{\text{SA}}}^{{}^{6}}$ unfeasible on a standard laptop. The set $\mathscr{R}_{{}_{\text{SA}}}^{{}^{7}}$ is certainly beyond reach, and the development of an algorithm that allows us to circumvent at least some of these difficulties will be the main goal of this paper.
@@
A simple observation, which is key to the development of this algorithm, is that already for small values of ${\sf N}$ the vast majority of extreme rays of the SACN violate SSA, i.e., $|\mathscr{R}_{{}_{\text{SSA}}}^{{}^{{\sf N}}}|\ll|\mathscr{R}_{{}_{\text{SA}}}^{{}^{{\sf N}}}|$.151515 This is evident already at ${\sf N}=3,4$ (e.g., see He:2022bmi ). Since we are only interested in $\mathscr{R}_{{}_{\text{SSA}}}^{{}^{{\sf N}}}$, one may then wonder whether there is a way to find this set more directly, without first deriving $\mathscr{R}_{{}_{\text{SA}}}^{{}^{{\sf N}}}$. A possible strategy would be to first find all the extreme rays of the “SSA-cone”, the polyhedral cone in entropy space carved out by all instances of SA and SSA. It would then appear straightforward to extract the subset of extreme rays which are also elements of $\mathscr{R}_{{}_{\text{SA}}}^{{}^{{\sf N}}}$. This strategy unfortunately suffers from the same issues described above, since the dimension of the ambient space is the same, the number of facets is comparable, and most extreme rays of this cone are not elements of $\mathscr{R}_{{}_{\text{SSA}}}^{{}^{{\sf N}}}$.
To focus on $\mathscr{R}_{{}_{\text{SSA}}}^{{}^{{\sf N}}}$ without having to introduce new inequalities, we will instead consider a variation of what is perhaps the most basic (and in general inefficient) strategy to find the extreme rays of any polyhedral cone. Consider a vector space $\mathbb{R}^{n}$ for some $n$, and a finite set $\mathcal{E}$ of linear inequalities that specify a full-dimensional pointed161616 A cone is pointed if it does not contain any linear subspace other than the trivial subspace. cone $\mathcal{C}$. The set of vectors that saturate an inequality is a hyperplane in $\mathbb{R}^{n}$. An arbitrary subset $\mathcal{X}\subseteq\mathcal{E}$ then determines a linear subspace of $\mathbb{R}^{n}$ given by the intersection of the hyperplanes in $\mathcal{X}$. If this subspace is $1$-dimensional, one can then check if it contains a ray that satisfies all the inequalities in $\mathcal{E}$. If such a ray exists, it is obviously an extreme ray of $\mathcal{C}$. By considering all possible subsets $\mathcal{X}\in 2^{\mathcal{E}}$, we can in principle find all the extreme rays. We call this strategy the “brute force strategy.”
@@
The structure of the paper is as follows. In §2, we review the necessary definitions about the mutual information poset, how it can be used to formulate KC, and a few additional results from our previous work that we will use to simplify the computation of $\mathscr{R}_{{}_{\text{SSA}}}^{6}$. The algorithm that we use is then presented in §3. We stress that this section is written independently from the set-up of the SAC reviewed in §2, and thus our algorithm can be immediately utilized for other applications. To exemplify how the algorithm works, we then re-compute the set $\mathscr{R}_{{}_{\text{SSA}}}^{5}$ in §4. The result of the computation for ${\sf N}=6$ is presented in §5, along with a preliminary analysis of this data and all the holographic graph models that we were able to construct. Readers who are only interested in the results of the ${\sf N}=6$ case, and not in the details of the algorithm, can skip earlier sections and proceed directly to §5.171717 Further details regarding the various attributes of $\mathscr{R}_{{}_{\text{SSA}}}^{6}$ are also presented in notes . Finally, in §6 we comment on open questions and future directions of investigation.

### v3

@@ 1 Introduction
The study of correlations is of central importance in quantum information theory, as it provides a comprehensive understanding of how information is distributed among different parts of a quantum system. Correlations are often explored through various measures and related constraints, which help in understanding the fundamental limits and capabilities of quantum systems, guiding the development of quantum technologies, and deepening our theoretical understanding of quantum mechanics. This work will focus on constraints which take the form of linear inequalities involving the von Neumann entropy of subsystems of a multipartite quantum system.
While some of these inequalities are “universal”, in the sense that they are obeyed by arbitrary quantum states, others are specific to particular classes of states, like stabilizers states Linden:2013kal ; gross-walter , or classical probability distributions 641556 ; 1998:Zhang ; 2002:Makarychev ; 2003:Zhang ; 2007:Matus . These “class dependent” inequalities therefore serve as characterizations of the states in these classes.
A class which is of particular interest for this work is that of “geometric states” in quantum gravity, or more specifically, in the context of gauge/gravity duality Maldacena:1997re ; Gubser:1998bc ; Witten:1998qj . These are the states in holographic CFTs which correspond to bulk classical geometries, for which the entropies of boundary spatial subsystems can be computed via the HRRT prescription Ryu:2006bv ; Hubeny:2007xt . The fact that these states obey entropy inequalities that are not universal was first noticed in Hayden:2011ag , and then explored more systematically in Bao:2015bfa . Since then, many new inequalities have been proven (for instance, see Czech:2022fzb ; Hernandez-Cuenca:2023iqh ; Czech:2024rco ) using the “contraction map” method of Bao:2015bfa , but a deeper conceptual understanding of the structure of the resulting holographic entropy cone (HEC) Bao:2015bfa for an arbitrary number of parties is still lacking. It is reasonable to expect that such an understanding would provide new insights about the entanglement structure of geometric states, and ultimately shed new light on how the bulk geometry is “encoded” in the boundary theory.
Motivated by this problem, and based on the ideas from Hubeny:2018trv ; Hubeny:2018ijt , a deep question about fundamental possible “patterns of dependences” in quantum mechanics was raised in Hernandez-Cuenca:2019jpv , which dubbed it the quantum marginal independence problem (QMIP). In its simplest and most intuitive form, the QMIP asks the following question: Given the binary specification of presence versus absence of correlations between all pairs of disjoint subsystems, does there exist any quantum state realizing such a choice? While this problem was formulated in Hernandez-Cuenca:2019jpv for arbitrary quantum states, and it is seemingly unrelated to questions concerning the structure of the HEC, it was argued in Hernandez-Cuenca:2022pst that the solution to the QMIP within the restricted class of geometric states (the holographic marginal independence problem (HMIP)) would provide the “essential data” from which the HEC can be reconstructed. In fact, Hernandez-Cuenca:2022pst even argued that it would be sufficient to only know the answer to the “extremal” version of the HMIP, where one only asks what are the possible situations where the number of independent subsystems is so large that demanding any additional independence would force the state to completely factorize. Formally, as we will review below, answering this question requires one to determine which extreme rays of the “subadditivity cone” can be realized by geometric states, or in practical terms, by the graph models of Bao:2015bfa .
Since the HEC is known explicitly for any number of parties ${\sf N}\leq 5$ Bao:2015bfa ; Cuenca:2019uzx , the conjectures in Hernandez-Cuenca:2022pst can easily be verified to be true in this context. However, while this is already a rather non-trivial check, it was previously observed in He:2023cco (using preliminary data from the present work) that the ${\sf N}=6$ case might have a much richer structure, and could therefore not only serve to provide additional highly non-trivial checks, but also to obtain more data that could guide intuition towards a characterization of the extremal marginal independence patterns which are holographic. With this as our main motivation, we develop in this work an algorithm to compute all the extreme rays of the ${\sf N}=6$ subadditivity cone that can possibly be realized by quantum states (not just geometric states in holography), or more precisely, all extreme rays that satisfy strong subadditivity. Below, we will review the main definitions and terminology, comment more technically on our motivations beyond the holographic framework, and outline the strategy that we will employ for the computation.
@@
In the holographic context, it was argued in Hernandez-Cuenca:2022pst that, for any given number of parties ${\sf N}$, if one knows the solution to the HMIP for each ${\sf N}^{\prime}\geq{\sf N}$, then one can in principle reconstruct all the inequalities that specify the ${\sf N}$-party holographic entropy cone. In fact, as we mentioned above, Hernandez-Cuenca:2022pst suggested that there might be an even more remarkable structure which only requires the solution to the extremal version of the HMIP rather than the full one. More precisely, let $\mathscr{R}_{{}_{\text{SA}}}^{{}^{{\sf N}}}$ be the set of all $1$-dimensional faces, or extreme rays (ERs), of the SACN, and $\mathscr{R}_{{}_{\text{Hol}}}^{{}^{{\sf N}}}\subseteq\mathscr{R}_{{}_{\text{SA}}}^{{}^{{\sf N}}}$ be the subset of these that can be realized holographically.555 Similarly, we will use $\mathscr{R}^{{}^{{\sf N}}}$ with other subscripts to indicate subsets of the SACN extreme rays pertaining to the class indicated by the subscript; refer to Table 1 for a summary of these. Then Hernandez-Cuenca:2022pst argued that for any ${\sf N}$, there exists some ${\sf N}^{\prime}\geq{\sf N}$ (which depends on ${\sf N}$) such that $\mathscr{R}_{{}_{\text{Hol}}}^{{}^{{\sf N}^{\prime}}}$ provides sufficient information to derive the extreme rays of the ${\sf N}$-party holographic entropy cone via a simple coarse-graining of subsystems. These then allow for a straightforward reconstruction of all the ${\sf N}$-party holographic entropy inequalities (although in practice this might be a hard computation). The result of Hernandez-Cuenca:2022pst is quite striking since the SAC itself is a much more primitive construct than the holographic entropy cone. It originates from the self-evident statement that the amount of correlation between any pair of subsystems cannot be negative, and as such it is not cognizant of holography. Indeed, the holographic input enters through the restriction to $\mathscr{R}_{{}_{\text{Hol}}}^{{}^{{\sf N}}}$.
@@
Interestingly, even though Hernandez-Cuenca:2022pst relied on technology only applicable in the holographic setting, a clue hinting at a similar structure in more general contexts was found recently in He:2023aif , which considered how to derive inner bounds to the general quantum entropy cone 1193790 ; 2005:Linden ; 2011:Cadney ; Christandl_2023 from the solution to the QMIP. Denoting by $\mathscr{R}_{{}_{\text{Sta}}}^{{}^{{\sf N}}}\subseteq\mathscr{R}_{{}_{\text{SA}}}^{{}^{{\sf N}}}$ the set of extreme rays of the SACN that can be realized by stabilizer states, the main result of He:2023aif was that a subset of $\mathscr{R}_{{}_{\text{Sta}}}^{{}^{9}}$ is sufficient to derive666 As in the holographic case, by derive we mean via subsystems coarse-grainings. an inner bound to the $4$-party quantum entropy cone which coincides with the full 4-party stabilizer cone Linden:2013kal .
@@
These results suggest that a similar structure might also exist at a more abstract level. For ${\sf N}\geq 3$, most faces of the SACN cannot be part of the solution to the QMIP for the simple reason that they are “ruled out” by strong subadditivity (SSA). Following He:2022bmi , we will say that a face is SSA-compatible if there is at least one entropy vector in its interior that satisfies all instances of SSA. For a fixed ${\sf N}$, the set of SSA-compatible faces has the structure of a lattice,777 A lattice is a poset $\mathcal{P}$ endowed with the additional property that for any two elements $x,y\in\mathcal{P}$, there exists both a least upper bound (called the join and denoted $x\vee y$) and a greatest lower bound (called the meet and denoted $x\wedge y$). where the partial order is given by reverse inclusion ($\mathscr{F}_{1}<\mathscr{F}_{2}$ if $\mathscr{F}_{1}\supset\mathscr{F}_{2}$), and the meet of two faces is the lowest dimensional face $\mathscr{F}=\mathscr{F}_{1}\wedge\mathscr{F}_{2}$ that includes $\mathscr{F}_{1}$ and $\mathscr{F}_{2}$ on its boundary He:2022bmi .888 The set of faces of an arbitrary polyhedral cone forms a lattice with the same partial order and meet operation. In general, the subset $\mathscr{R}_{{}_{\text{SSA}}}^{{}^{{\sf N}}}\subseteq\mathscr{R}_{{}_{\text{SA}}}^{{}^{{\sf N}}}$ of SSA-compatible extreme rays of the SACN is a subset of the set of coatoms of this lattice,999 For ${\sf N}\leq 4$, each coatom is an extreme ray, but it is not known if the inclusion is strict for larger ${\sf N}$; see He:2022bmi for more details. but He:2022bmi showed that for ${\sf N}\geq 4$ the lattice is not coatomistic. This means that not every element of the lattice can be obtained as the meet of a collection of coatoms. In other words, even geometrically, knowing $\mathscr{R}_{{}_{\text{SSA}}}^{{}^{{\sf N}}}$ is in general insufficient information to be able to recover all the SSA-compatible higher-dimensional faces.101010 Technically, this result was proven for the set of SSA-compatible faces, which are not necessarily all realizable by quantum states. Therefore, it could have been in principle possible that each element of the lattice which is not the meet of a collection of coatoms is non-realizable, implying that the sublattice of realizable elements is coatomistic. However, we will see momentarily that already for ${\sf N}=4$ this is not true. This result strongly suggests that in general, the solution to the QMIP for fixed ${\sf N}$ cannot be deduced from $\mathscr{R}_{{}_{\text{QM}}}^{{}^{{\sf N}}}$, which is only the solution to the extremal version of QMIP involving the same ${\sf N}$, and that one needs to know all the meet-irreducible elements of the lattice.111111 An element $x$ of a lattice is said to be meet-irreducible if $x=y\wedge z$ implies $x=y$ or $x=z$. For ${\sf N}=4$ however, it was already shown in Hernandez-Cuenca:2019jpv that all SSA-compatible faces of the SAC4 are realizable in quantum mechanics, and that in fact stabilizer states are sufficient. This implies that the meet-irreducible elements of the lattice of SSA-compatible faces of the SAC4 which are not coatoms can be obtained from coarse-grainings of certain coatoms of the analogous lattice for ${\sf N}=9$. In other words, similar to the holographic case, the solution to the QMIP can be derived from the solution to its extremal version involving a larger number of parties. All these observations suggest that there might be classes of quantum states for which the solution to the extremal marginal independence problem for all ${\sf N}^{\prime}\geq{\sf N}$ carries a lot of (if not complete) information about the solution to the (seemingly) more general ${\sf N}$-party QMIP, and may even contain information about entropy inequalities. This motivates us to study the set $\mathscr{R}_{{}_{\text{SSA}}}^{{}^{{\sf N}}}$ more closely.
@@
Another important motivation for us to study $\mathscr{R}_{{}_{\text{SSA}}}^{{}^{{\sf N}}}$ comes again from holography. Suppose that the conjecture proposed in Hernandez-Cuenca:2022pst is true, and that all ${\sf N}$-party holographic entropy inequalities can indeed be recovered from $\mathscr{R}_{{}_{\text{Hol}}}^{{}^{{\sf N}^{\prime}}}$ for some ${\sf N}^{\prime}\geq{\sf N}$. Then what characterizes the extreme rays that can be realized in quantum mechanics but not by geometric states? A first guess, informed by data for ${\sf N}\leq 5$, is that no such rays exist and $\mathscr{R}_{{}_{\text{QM}}}^{{}^{{\sf N}}}=\mathscr{R}_{{}_{\text{Hol}}}^{{}^{{\sf N}}}$. In other words, all extreme rays of the SACN that can be realized in quantum mechanics would also be realizable holographically. However, this was shown to be false in He:2023cco , which proved the strict inclusion $\mathscr{R}_{{}_{\text{Hol}}}^{{}^{{\sf N}}}\subset\mathscr{R}_{{}_{\text{QM}}}^{{}^{{\sf N}}}$ for ${\sf N}\geq 6$ via the construction of an explicit counterexample, namely an extreme ray of the SAC6 that can be realized by a stabilizer state but not by a geometric state.
@@
where {${n\brace k}$ is the Stirling number of the second kind Hubeny:2018ijt .121212 The number of SA instances is {${{\sf N}+2\brace 3}$, but the $12^{{\sf N}}-$ 1 instances of the form ${\sf S}_{\mathscr{J}}\geq 0$ are not facets He:2022bmi . For ${\sf N}\leq 4$ the problem is easily manageable by freely available software Normaliz:301 , but the ${\sf N}=5$ case is already significantly more involved, since ${\sf{D}}=31$ and there are $270$ facets. The computation can be largely simplified using a result of He:2022bmi , which identified for arbitrary ${\sf N}$ a specific lower-dimensional face $\mathscr{F}_{*}$ of the SACN that includes all elements of $\mathscr{R}_{{}_{\text{SSA}}}^{{}^{{\sf N}}}$ (with the exception of the trivial ones corresponding to the Bell pairs).131313 Since we will utilize this result in the computation presented in this paper, we will review it in detail in the next section. In general, the dimension of $\mathscr{F}_{*}$ is given by
@@
which for ${\sf N}=5$ gives $\text{dim}(\mathscr{F}_{*})=16$. Furthermore, upon dimensional reduction of the inequalities that specify the SAC5 to the linear subspace spanned by $\mathscr{F}_{*}$, many inequalities become redundant, and $\mathscr{F}_{*}$ only has $80$ facets. In this manner, the computation of $\mathscr{R}_{{}_{\text{SA}}}^{{}^{5}}$ becomes straightforward using Normaliz:301 , but as one can immediately see from (4) and (5), this simplification quickly becomes insufficient as ${\sf N}$ grows. Indeed, already at ${\sf N}=6$, the face $\mathscr{F}_{*}$ has $\text{dim}(\mathscr{F}_{*})=42$ and is specified by $280$ inequalities (the number of facets of SAC6 is $903$), rendering the computation of $\mathscr{R}_{{}_{\text{SA}}}^{{}^{6}}$ unfeasible on a standard laptop. The set $\mathscr{R}_{{}_{\text{SA}}}^{{}^{7}}$ is certainly beyond reach, and the development of an algorithm that allows us to circumvent at least some of these difficulties will be the main goal of this paper.
@@
A simple observation, which is key to the development of this algorithm, is that already for small values of ${\sf N}$ the vast majority of extreme rays of the SACN violate SSA, i.e., $|\mathscr{R}_{{}_{\text{SSA}}}^{{}^{{\sf N}}}|\ll|\mathscr{R}_{{}_{\text{SA}}}^{{}^{{\sf N}}}|$.141414 This is evident already at ${\sf N}=3,4$ (e.g., see He:2022bmi ). Since we are only interested in $\mathscr{R}_{{}_{\text{SSA}}}^{{}^{{\sf N}}}$, one may then wonder whether there is a way to find this set more directly, without first deriving $\mathscr{R}_{{}_{\text{SA}}}^{{}^{{\sf N}}}$. A possible strategy would be to first find all the extreme rays of the “SSA-cone”, the polyhedral cone in entropy space carved out by all instances of SA and SSA. It would then appear straightforward to extract the subset of extreme rays which are also elements of $\mathscr{R}_{{}_{\text{SA}}}^{{}^{{\sf N}}}$. This strategy unfortunately suffers from the same issues described above, since the dimension of the ambient space is the same, the number of facets is comparable, and most extreme rays of this cone are not elements of $\mathscr{R}_{{}_{\text{SSA}}}^{{}^{{\sf N}}}$.
To focus on $\mathscr{R}_{{}_{\text{SSA}}}^{{}^{{\sf N}}}$ without having to introduce new inequalities, we will instead consider a variation of what is perhaps the most basic (and in general inefficient) strategy to find the extreme rays of any polyhedral cone. Consider a vector space $\mathbb{R}^{n}$ for some $n$, and a finite set $\mathcal{E}$ of linear inequalities that specify a full-dimensional pointed151515 A cone is pointed if it does not contain any linear subspace other than the trivial subspace. cone $\mathcal{C}$. The set of vectors that saturate an inequality is a hyperplane in $\mathbb{R}^{n}$. An arbitrary subset $\mathcal{X}\subseteq\mathcal{E}$ then determines a linear subspace of $\mathbb{R}^{n}$ given by the intersection of the hyperplanes in $\mathcal{X}$. If this subspace is $1$-dimensional, one can then check if it contains a ray that satisfies all the inequalities in $\mathcal{E}$. If such a ray exists, it is obviously an extreme ray of $\mathcal{C}$. By considering all possible subsets $\mathcal{X}\in 2^{\mathcal{E}}$, we can in principle find all the extreme rays. We call this strategy the “brute force strategy.”
@@
The structure of the paper is as follows. In §2, we review the necessary definitions about the mutual information poset, how it can be used to formulate KC, and a few additional results from our previous work that we will use to simplify the computation of $\mathscr{R}_{{}_{\text{SSA}}}^{6}$. The algorithm that we use is then presented in §3. We stress that this section is written independently from the set-up of the SAC reviewed in §2, and thus our algorithm can be immediately utilized for other applications. To exemplify how the algorithm works, we then re-compute the set $\mathscr{R}_{{}_{\text{SSA}}}^{5}$ in §4. The result of the computation for ${\sf N}=6$ is presented in §5, along with a preliminary analysis of this data and all the holographic graph models that we were able to construct.161616 Further details regarding the various attributes of $\mathscr{R}_{{}_{\text{SSA}}}^{6}$ are also presented in rays-notes , and the actual graph constructions in graph-constructions . Readers who are only interested in the results of the ${\sf N}=6$ case, and not in the details of the algorithm, can skip earlier sections and proceed directly to §5. Finally, in §6 we comment on open questions and future directions of investigation.